# Quantum-thermodynamic enhancements in continuous thermal machines require energetic coherence

### v1

@@ VII Conclusions
We have characterized the thermodynamic impact of energetic and degenerate coherence in the performance of quantum thermal machines operating in nonequilibrium steady states. While machines displaying energetic coherence in their steady states can lead to quantum-thermodynamic advantages in terms of their tradeoff between power, efficiency and stability, this is not the case for machines only displaying coherence between degenerate levels, no matter its origin being Hamiltonian or noise-induced. These results, obtained through direct comparison of prototypical quantum thermal machines with axiomatically-constructed classical equivalent models, imply that many previous claims regarding quantum-enhancements in these machines are spurious. Indeed, we showed that even if the dynamics of these machines exhibit purely quantum features (such as e.g. entanglement), a classical Markovian machine using the same set of resources (energy gaps and bath temperatures) can be systematically constructed that performs just as well or better than the original machine. Our results provide, at the same time, an explanation of why violations of the TUR cannot be found in devices showing degenerate coherence.
@@
Using the three-level amplifier as a main example of a quantum thermal machine displaying energetic coherence, we have shown that it always match or exceeds the performance of its classical equivalent counterpart in all possible configurations, with improvements in the machine stability that become greater for larger temperature bias, i.e. far from equilibrium. The parameter regions where these improvements are maximal coincide indeed with regimes where the thermal machine breaks the TUR bound. Moreover, we also observe wide regions showing (smaller) thermodynamic improvements that are not witnessed by TUR violations. Importantly, our results reveal that in order to observe a truly quantum-thermodynamic advantage in this model, it is necessary (and sufficient) to consider the fluctuations in the currents (at least at the level of the variance), in contrast to previous assessments based on more limited notions of classical equivalents (leading to a less stringent comparison) [9].
Beyond the direct comparison between quantum machines and their respective equivalents for given parameters, we also provided an all-to-all comparison computing the optimal configurations (Pareto front) of quantum and classical machines maximizing power and efficiency while minimizing fluctuations. Our analysis reveals that quantum-thermodynamic advantages in the three-level amplifier occur indeed within the optimal front, allowing thus for a practical quantum-thermodynamic advantage. Since implementations of this or similar thermal machine models have been proposed in a number of platforms [81, 13], and it has been actually experimentally realized with NV-centers in diamond [12], it would be interesting to test our results for the quantum-thermodynamic advantages reported here in relevant regimes by suitably measuring the variance of the currents [82, 83].
@@ Appendix A General expression of current first moments with local dissipation
$\begin{split}&\langle\dot{Q}_{r}\rangle=\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]]=\sum_{k}\operatorname{Tr}[\mathcal{D}_{k}^{\dagger(r)}[H_{0}]\pi(t)]\\ &=\sum_{k}\sum_{m,l}\bra{m}L_{k}^{(r)\dagger}H_{0}L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},H_{0}\}\ket{l}\pi_{lm}\end{split}$ (37)
where we have applied the cyclic property of the trace to obtain the adjoint dissipator $\mathcal{D}_{k}^{\dagger(r)}[\cdot]:=L_{k}^{(r)\dagger}\cdot L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},\cdot\}$. Now, using the explicit form of the Lindblad operators in Eq. (4) and the bare Hamiltonian, $H_{0}=\sum_{i}\epsilon_{i}\ket{i}\bra{i}$ |, we can obtain each term in the sum of the last expression:
@@ Appendix B Full Counting Statistics
The Full Counting Statistics (FCS) formalism is used to compute the variances of the different input and output currents of the quantum thermal machines presented and their classical equivalent models. In this formalism a set of counting fields $\{\chi_{r}\}$ are introduced that keep track of the exchanges of energy quanta between the machine and the thermal reservoirs $r$. These lead to the derivation of a generalized master equation for an extended density operator $\rho_{G}(t,\{\chi_{k}^{r}\}$ ) depending on the fields. It reads [78]:
@@
$\begin{split}\bar{\mathcal{D}}^{(r)}_{k}\left[\rho_{\mathrm{G}}\right]:=\left(e^{-\nu_{k}^{(r)}\chi_{r}}~{}L_{k}^{(r)}\rho_{\mathrm{G}}L_{k}^{(r)~{}\dagger}-\dfrac{1}{2}\left\{L_{k}^{(r)\dagger}L^{(r)}_{k},\rho_{\mathrm{G}}\right\}\right).\end{split}$ (43)
@@
In any case, in the limit $\{\chi_{r}\}\rightarrow 0$, we recover $\rho_{\mathrm{G}}(t)=\rho(t)$ and (B) reduces to the standard master equation (2) for the machine evolution. Moreover, as in the case of the original master equation, Eq. (B) can be linearised and written in the form
@@
for $n=1,2,3,...$. Here the first ($n=1$) and second ($n=2$) cumulants correspond, respectively, to the average ($\mathcal{C}_{1}^{(r)}=\langle\dot{N}_{r}\rangle$) and variance ($\mathcal{C}_{2}^{(r)}=\mathrm{Var}[\dot{N}_{r}]$) of the currents of quanta on that reservoirs. The average and variances of the heat currents in which we are mainly interested in this work are then given by:
@@
Analogously, the average and variance of the power for the case of the three-level amplifier are given, respectively, by the first law, $\langle\dot{W}\rangle=\sum_{r}\langle\dot{Q}_{r}\rangle$ and in the long-time limit, $\mathrm{Var}[\dot{W}]=\sum_{r}\mathrm{Var}[\dot{Q}_{r}]+2\mathrm{cov}[\dot{Q}_{\mathrm{c}}\dot{Q}_{\mathrm{h}}]=\omega_{d}^{2}\mathrm{Var}[\dot{N}]$ for $r=\mathrm{c,h}$ where the last equality follows from $\mathrm{Var}[\dot{Q}_{r}]=\Delta\epsilon_{r}^{2}\mathrm{Var}[\dot{N}]$ and $\mathrm{cov}[\dot{Q}_{\mathrm{c}}\dot{Q}_{\mathrm{h}}]=\Delta\epsilon_{\mathrm{h}}\Delta\epsilon_{\mathrm{c}}\mathrm{Var}[\dot{N}]$.
Unfortunately the size and complexity of the matrix $W_{G}$ makes in many cases impossible to obtain analytically the eigenvalue $\lambda(\{\chi_{r}\})$ by direct diagonalization of $W_{G}$, and other methods are required. In order to obtain the first and second cumulants analytically, we follow the method known as “inverse full counting” originally introduced in Ref. [79] and used in Refs. [30, 33] for similar purposes. In the following we review this method for the case of a single field $\chi$, but the expressions can be naturally extended to multiple fields $\{\chi_{r}\}$ as it is our case here (see e.g. appendix C in Ref. [33]).
@@
$\mathcal{C}_{2}=\dfrac{2(a_{0}^{\prime}a_{1}a_{1}^{\prime}-a_{0}^{\prime 2}a_{1}^{\prime})-a_{0}^{\prime\prime}a_{1}^{2}}{a_{1}^{3}}.$ (55)
In the following we provide the form of the matrix $W_{G}$ for the three quantum thermal machines used as main illustrative examples in this paper and their corresponding classical equivalents. The corresponding expressions for the cumulants are, however, not included here for size reasons.
B.1 Three-level coherent amplifier
associated to vector $\vec{p}_{G}(t)=(\rho_{00},\rho_{11},\rho_{22},\mathrm{Im}[\rho_{01}])$. In the case of the classical equivalent, we no longer consider the contribution of coherence to the dynamics, but introduce the classical stochastic transition rate between the interacting levels, leading to a matrix $W_{G}^{\mathrm{cl}}$ that reads:
with the corresponding associated vector $\vec{p}_{G}^{\,\mathrm{cl}}(t)=(\rho_{00},\rho_{11},\rho_{22})$.
B.2 Three-qubit autonomous refrigerator
associated to vector $\vec{p}_{G}(t)=(\rho_{000},\rho_{001},\rho_{100},\rho_{011},\rho_{111},\rho_{101},\rho_{010},\mathrm{Im}[\rho_{010-101}])$. For the classical equivalent we instead loose the coherence contribution, while adding the extra rate $\gamma^{\mathrm{cl}}$ between degenerated levels:
for vector $\vec{p}_{G}^{\,\mathrm{cl}}(t)=(\rho_{000},\rho_{001},\rho_{100},\rho_{011},\rho_{111},\rho_{101},\rho_{010})$. In the above equations, we defined $\Gamma_{1}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{2}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{3}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{4}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{5}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{6}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{7}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{8}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$ and $\Gamma_{9}=-\frac{1}{2}(\gamma_{d}+\gamma_{2d}+\gamma_{3d}+\gamma_{u}+\gamma_{2u}+\gamma_{3u})$.
B.3 Noise-induced-coherent machine
with $\vec{p}_{G}(t)=(\rho_{00},\rho_{11},\rho_{\alpha\alpha},\rho_{\beta\beta},\mathrm{Re}[\rho_{\alpha\beta}])$. The matrix for the classical equivalent of the NIC machine does not contain the coherence anymore, but the modified rates $\{\gamma_{1\alpha}^{\mathrm{cl}},\gamma_{\alpha 1}^{\mathrm{cl}},\gamma_{1\beta}^{\mathrm{cl}},\gamma_{\beta 1}^{\mathrm{cl}},\gamma_{\alpha\beta}^{\mathrm{cl}}\}$. It reads:
with associated vector $\vec{p}_{G}^{\,\mathrm{cl}}(t)=(\rho_{00},\rho_{11},\rho_{\alpha\alpha},\rho_{\beta\beta})$.
@@ Appendix C Classical thermodynamic-equivalent models details
for $n=\{\text{u},\text{v}\}$. They result in a steady-state coherence term given by (28) where the introduction of (28) into (62) leads to corrections for rates and a new transition between levels $\mathrm{u}$ and $\mathrm{v}$ as shown in (29).

### v2

@@ VIII Conclusions
We have characterized the thermodynamic impact of Hamiltonian and noise-induced coherence in the performance of quantum thermal machines operating in nonequilibrium steady states. While machines displaying Hamiltonian coherence in their steady states lead to quantum-thermodynamic advantages in terms of their tradeoff between power, efficiency and stability, this is not always the case for machines displaying noise-induced coherence. These results, obtained through direct comparison of generic quantum thermal machines with axiomatically-constructed classical equivalent models, ensure the presence of quantum-enhanced performance for models of Hamiltonian-induced coherence in out of equilibrium situations. On the other hand, they also imply that several previous claims of thermodynamic advantages in machines with noise-induced coherence may be spurious, and suggests that they should be reassessed on a case-by- case basis, by determining the set of parameters in which quantum advantages may really exist (if any). Indeed, even if the dynamics of these machines may exhibit purely quantum features, a classical Markovian machine using the same set of resources (energy gaps and bath temperatures) can be systematically constructed that performs just as well or better than the original machine.
@@
Using the three-level amplifier and the three-qubit-fridge as main examples of a quantum thermal machines displaying Hamiltonian coherence (the first energetic coherence and the second degenerate one), we have illustrated situations where coherence is always beneficial. In these cases the performance match or exceeds the performance of its classical equivalent counterpart in all possible configurations, with improvements in the machine stability that become greater far from equilibrium. In the case of the three-level amplifier, the parameter regions where these improvements are maximal coincide indeed with regimes where the thermal machine breaks the TUR bound. However, we also observe wide regions showing (smaller) thermodynamic improvements that are not witnessed by TUR violations. This is indeed entirely the case for the three-qubit-fridge, where the TUR by itself is not able to spot thermodynamic enhancements. Importantly, our results reveal that in order to observe a genuine quantum-thermodynamic advantage, it is necessary (and sufficient) to consider the fluctuations in the currents (at least at the level of the variance), in contrast to previous assessments based on more limited notions of classical equivalents (leading also to a less stringent comparison) [9].
Beyond the direct comparison between quantum machines and their respective equivalents for given parameters, we also provided an all-to-all comparison computing the optimal configurations (Pareto front) of quantum and classical machines maximizing power and efficiency while minimizing fluctuations. Our analysis reveals that quantum-thermodynamic advantages in both the three-level amplifier and the three-qubit-fridge occur indeed within the optimal front, allowing thus for a practical quantum-thermodynamic advantage. However, it is also clear from our analysis that in the three-qubit-fridge, the advantages are much lower than in the three-level amplifier. Since implementations of these or similar thermal machine models have been proposed in a number of platforms [58, 87, 57, 13, 56], and have been actually experimentally realized with NV-centers in diamond [12] and trapped ions [11] respectively, it would be interesting to test our results for the quantum-thermodynamic advantages reported here in relevant regimes by suitably measuring the variance of the currents [94, 95].
@@ Appendix A General expression of current first moments with local dissipation
$\begin{split}&\langle\dot{Q}_{r}\rangle=\sum_{k}\operatorname{Tr}[(H_{0}+V(t))\mathcal{D}_{k}^{(r)}[\pi(t)]]\\ &=\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]]=\sum_{k}\operatorname{Tr}[\mathcal{D}_{k}^{\dagger(r)}[H_{0}]\pi(t)]\\ &=\sum_{k}\sum_{m,l}\bra{m}L_{k}^{(r)\dagger}H_{0}L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},H_{0}\}\ket{l}\pi_{lm}\end{split}$ (43)
where the terms $\operatorname{Tr}[V(t)\mathcal{D}_{k}^{(r)}[\pi(t)]]\sim g\gamma_{0}^{k}$ have been neglected in accordance to the weak coupling and weak driving approximations. We have applied the cyclic property of the trace to obtain the adjoint dissipator $\mathcal{D}_{k}^{\dagger(r)}[\cdot]:=L_{k}^{(r)\dagger}\cdot L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},\cdot\}$. Now, using the explicit form of the Lindblad operators in Eq. (4) and the bare Hamiltonian, $H_{0}=\sum_{i}\epsilon_{i}\ket{i}\bra{i}$ |, we can obtain each term in the sum of the last expression:
@@ Appendix B Full Counting Statistics
The Full Counting Statistics (FCS) formalism is used to compute the variances of the different input and output currents of the quantum thermal machines presented and their classical equivalent models. In this formalism a set of counting fields $\vec{\chi}:=\{\chi_{r}\}$ are introduced that keep track of the exchanges of energy quanta between the machine and the thermal reservoirs $r$. These lead to the derivation of a generalized master equation for an extended density operator $\rho_{G}(t,\{\chi_{r}\}$ ) depending on the fields. It reads [78]:
@@
$\begin{split}\bar{\mathcal{D}}^{(r)}_{k}\left[\rho_{\mathrm{G}}\right]:=&e^{-\nu_{k}^{(r)}\chi_{r}}~{}L_{k}^{(r)}\rho_{\mathrm{G}}L_{k}^{(r)~{}\dagger}\\ &-\dfrac{1}{2}\left\{L_{k}^{(r)\dagger}L^{(r)}_{k},\rho_{\mathrm{G}}\right\}.\end{split}$ (49)
@@
In any case, in the limit $\{\chi_{r}\}\rightarrow 0$, we recover $\rho_{\mathrm{G}}(t)=\rho(t)$ and (B) reduces to the standard master equation (2) for the machine evolution. Moreover, as in the case of the original master equation, Eq. (B) can be linearized and written in the form
@@
for $n=1,2,3,...$. Here the first ($n=1$) and second ($n=2$) cumulants correspond, respectively, to the average ($\mathcal{C}_{1}^{(r)}=\langle\dot{N}_{r}\rangle$) and variance ($\mathcal{C}_{2}^{(r)}=\mathrm{Var}[\dot{N}_{r}]$) of the currents of quanta on that reservoir. The average and variances of the heat currents in which we are mainly interested in this work are then given by:
@@
Due to the preservation of quanta exchanges in all the interactions in the weak coupling limit, it follows that for long trajectories in steady-state machines where the different reservoirs contribute the same number of quanta along all possible cycles, we have a proportionality among all currents. That means that the fluctuations of every current are also proportional:
$\begin{split}\mathrm{Var}[\dot{Q}_{r}]&=\Delta\epsilon_{r}^{2}\mathrm{Var}[\dot{N}],\\ \mathrm{Cov}[\dot{Q}_{r}\dot{Q}_{s}]&=\Delta\epsilon_{r}\Delta\epsilon_{s}\mathrm{Var}[\dot{N}],\end{split}$ (54)
where ${\rm Var}[\dot{N}]={\rm Var}[\dot{N}_{r}]$ for all $r$. Analogously, the average and variance of the power for the case of external driving are given, respectively, by the first law, $\langle\dot{W}\rangle=\sum_{r}\langle\dot{Q}_{r}\rangle$, and in the long-time limit we have:
$\begin{split}\mathrm{Var}[\dot{W}]&=\sum_{r}\left(\mathrm{Var}[\dot{Q}_{r}]+2\sum_{s}\mathrm{Cov}[\dot{Q}_{r}\dot{Q}_{s}]\right)\\ &=\omega_{d}^{2}\mathrm{Var}[\dot{N}].\end{split}$ (55)
where we have used Eqs. (54).
Unfortunately the size and complexity of the matrix $W_{G}$ in many cases makes impossible to obtain analytically the largest eigenvalue $\lambda(\{\chi_{r}\})$ by direct diagonalization of $W_{G}$, and other methods are required. In order to obtain the first and second cumulants analytically, we follow the method known as “Inverse Counting Statistics” originally introduced in Ref. [79] and used recently in Refs. [30, 33] for similar purposes. In the following we review this method for the case of a single field $\chi$, which is enough for obtaining all the relevant quantities in our case, but the expressions can be naturally extended to multiple fields $\{\chi_{r}\}$ (see e.g. Appendix C in Ref. [33]).
@@
$\mathcal{C}_{2}=-\dfrac{1}{a_{1}}\left(a_{0}^{\prime\prime}+a_{1}^{\prime}\mathcal{C}_{1}+a_{2}\mathcal{C}_{1}^{2}\right).$ (63)
Replacing the expression for the cumulant $\mathcal{C}_{1}$ in Eqs. 62 into the above Eq. (63), we obtain the final expression for the variance of the probability current in the long-time limit:
${\rm Var}[\dot{N}]=-\frac{1}{a_{1}}[a_{0}^{{}^{\prime\prime}}-a_{1}^{{}^{\prime}}\Big{(}\frac{a_{0}^{{}^{\prime}}}{a_{1}}\Big{)}+a_{2}\Big{(}\frac{a_{0}^{{}^{\prime}}}{a_{1}}\Big{)}^{2}],$ (64)
which is used to obtain all the variances for heat and work currents [Eqs. (54) and (55)]
@@ Appendix C Classical thermodynamic-equivalent models details
for $n=\{\text{u},\text{v}\}$. They result in a steady-state coherence term given by (28). The introduction of (28) into (71) leads to corrections for the rates and a new stochastic transition between levels $\mathrm{u}$ and $\mathrm{v}$ as shown in Eq. (29).